OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−5, −5) and a directrix of y = 7.

OpenStudy (anonymous):

f(x) = −one twenty-fourth (x − 1)2 − 5 f(x) = one twenty-fourth (x − 1)2 − 5 f(x) = −one twenty-fourth (x + 5)2 + 1 f(x) = one twenty-fourth (x + 5)2 + 1

OpenStudy (anonymous):

Someone is CHEATING smh!

OpenStudy (anonymous):

._. How is asking a question cheating? It was really nice of you to go through my profile, though. I've also already answered this question on my own, thanks.

OpenStudy (anonymous):

oh but asking if anyone takes a certain class is cheating though ? and how about you unblock me i have something to say to you in your inbox

OpenStudy (anonymous):

nothing mean -_-

OpenStudy (anonymous):

Yes, asking if someone is taking a specific course is cheating because we all know the next step you plan to take ._. And that is asking them for specific help on specific tests. When instead you can just post your answer and get help. I really feel you're lying so, no, not right now.

OpenStudy (anonymous):

If someone is in my class than they can help me not give me the answers why would i want someone who isn't in my class and they " THINK " the answer is B or something -_- and trust me im not lying believe me..

OpenStudy (anonymous):

Go for it. And because you can do the same thing by posting a question. Like I've done above.